A certain amount of metal whose equivalent mass is 28 displaces 0.7L of H2 at STP from an acid, hence the mass of the element is _?

Ans: 1.75g

Grade:Upto college level

3 Answers

Sunil Kumar FP
10 years ago
equvalent weight of metal will be equivalentnweight of H2
no of mole pf h2 formed =pv/rt =1×.7/273×.0821 =.0312
let the weight of metal be x
x/28=.0312/2
x=1.747 g

thanks and regards
sunil kr
askIItian faculty
Yashasvee
7 years ago
EQmetal=EQHydrogen 
If 22.4l of hydrogen have 2 gm or 1 mole 
So 0.7 have =0.7/22.4X2=0.06gm of H2.   Now let mass of metal be x then according to our first line eq metal =eq hydrogen     That is.  X/28=0.06/1=0.17
ABC
5 years ago
Equivalent w.t. of metal ÷ Equivalent w.t. of H2 = E/1-------(a)Also (mass of metal/ mass of H2 displaced) = E/1 -------(b)given E = 28 We know that 1 mole of H2 i.e of 2gm have volume = 22.4L so in .7 L .06gm H2 is present So according to formula (b) mass of metal/ mass of H2 displaced = E/1So mass of metal/.06 =28/1mass of metal= 1.75 g ( app.)